1. Assuming an annual interest rate of 0:10 (10%), find the presented discounted value: a. of $1100 received at end of one year; plus $1210 received at end of two years; plus $1331 received at end of three years [hint: this is the sum of three numbers].b. $2420 received at the end of two years.
Answer #1

a) Presented discount value = 1100*(P/F,10%,1)+1210*(P/F,10%,2)+1331*(P/F,10%,3)

= 1100*0.909091+1210*0.826446+1331*0.751315

= 3000

b) Presented discount value = 2420*(P/F,10%,2)

= 2420*0.826446

= 2000